OpenStudy (anonymous):

Please help! True or false? If the center of a circle is (0,0) and the radius is sq.rt.7, the equation of the circle is x^2+y^2=7.

OpenStudy (anonymous):

true

OpenStudy (anonymous):

How come? Please explain :)

OpenStudy (anonymous):

the equation of the circle is given by (x-h)^2+(y-k)^2=r^2. Here,(h,k) is the centre and r is the radius.Plug the values (0,0) and sqrt(7) in the equation,you get, x^2+y^2=7